OpenStudy (anonymous):

Maria solved the equation below and got x = –4. Which of the following choices verifies that her solution is correct? 5x + 22 = x + 6 A. 5(–4) + 22 = (–4) + 6 –20 + 22 = –2 –2 = –2 B. 5(–31) + 22 = (–10) + 6 –26 + 22 = –4 –4 = –4 C. 5(–4) + 22 = (–4) + 6 5(18) = (–4) +6 24 = 24 D. 5(–4) + 22 = (–4) + 6 –20 + 22 = 2 2 = 2

OpenStudy (camerondoherty):

Well, lets see

OpenStudy (camerondoherty):

x=-4

OpenStudy (camerondoherty):

we know tht

OpenStudy (aum):

To verify the solution, replace x with -4 and make sure the LHS equals the RHS. LHS - Left Hand side

OpenStudy (camerondoherty):

so 5(-4) + 22 = (-4) + 6

OpenStudy (camerondoherty):

5(-4) + 22 = (-4) + 6 -20 + 22 = 2 2=2

OpenStudy (camerondoherty):

So which one would it be?

OpenStudy (anonymous):

d